# Small-time asymptotics and the emergence of complex singularities for the KdV equation

### v1

@@ B.3 Special case $A_{1}=0$
Very briefly, if $A_{1}=0$, then the solution for $f_{1}$ is simply $f_{1}=0$, so that (105) implies $\mu_{1}=0$. Further, the ode for $f_{2}$, (99), becomes
$\xi f_{2}^{\prime}+6(f_{0}f_{2}^{\prime}+f_{0}^{\prime}f_{2})+f_{2}^{\prime\prime\prime}=0,$ (115)
and the local behaviour near $\xi=\xi_{0}$, namely (106), simplifies to
$f_{2}\sim\frac{\mu_{2}}{(\xi-\xi_{0})^{3}}\quad\mbox{as}\quad\xi\rightarrow\xi_{0}.$ (116)
@@
where the value $\mu_{2}$ comes from solving (115) subject to
$f_{2}\sim A_{2}+\frac{4A_{0}A_{2}}{\xi^{3}}+\frac{40A_{0}A_{2}(A_{0}+2)}{\xi^{6}}\quad\mbox{as}\quad\xi\rightarrow-\mathrm{i}\infty$ (117)
@@
An interesting check on these arguments is with $A_{0}=-6$, $A_{1}=0$ and $A_{2}=2$, which is what we have for the initial condition $u_{0}=6\,\mathrm{sech}^{2}x$ and the corresponding $2$-soliton solution (76). In that case, we have
@@
as can be checked by substitution into (27) with (28) and (115) with (117) using $A_{0}=-6$ and $A_{2}=2$. In this exceptional case, there are only three singularities of $f_{0}$, which are given by (56). For each of these, the near-singularity behaviour for $f_{2}$ is
@@
which, by comparison with (116), means that $\mu_{2}=4$ and the three singularities emerge out of $x=x_{0}$ according to
$s(t)\sim x_{0}+(3t)^{1/3}\xi_{0}+4(3t)\quad\mbox{as}\quad t\rightarrow 0^{+}.$ (118)
@@
Note the special case $A_{0}=-12$, $A_{1}=0$, $A_{2}=4$ is relevant for the initial condition $u_{0}=12\,\mathrm{sech}^{2}x$, which leads to the $3$-soliton (80) described in subsection 4.3. In this case, the solutions of (27) with (28) and (115) with (117) are
@@
$s(t)\sim x_{0}+(3t)^{1/3}\xi_{0}+8(3t)\quad\mbox{as}\quad t\rightarrow 0^{+}.$ (119)
The asymptotic prediction (119) agrees with the exact results for the $3$-soliton solution in subsection 4.2.
@@ B.4 Special case $A_{0}=-2$
In terms of our original problem (1) with (3), the case $A_{0}=-2$ is subtly different because this is when the initial condition has a double pole that has the same strength exhibited by the double poles that time-dependent solutions to KdV have. We therefore expect the temporal scaling to be different for $A_{0}=-2$, with any singularity moving more slowly away from $x=x_{0}$ than those that are born for $A_{0}\neq-2$.
To help clarify these cases for $A_{0}=-2$, we observe that (107) means that singularities of solutions of KdV are doubles poles with zero residue. Thus we have two broad cases to consider. First, if $A_{0}=-2$ but $A_{1}\neq 0$, then the initial condition has the same leading-order behaviour near $x=x_{0}$ as time-dependent solutions of KdV have near each singularity $x=s(t)$, but with a different first-order correction. It turns out, in this case, that singularities are modified at $x=x_{0}$ but move with initial speed $\mathcal{O}(t^{-1/3})$ (instead of $\mathcal{O}(t^{-2/3})$, which is what happens if $A_{0}\neq-2$). Second, if $A_{0}=-2$ and $A_{1}=0$, then the initial condition has the same leading-order and correction term near its singularity as the full time-dependent solution does. This would happen, of course, if we take any time-dependent solution of (1) (that has evolved for any nonzero time) and treat it as an initial condition. Here the singularity present at $x_{0}$ propagates with speed $\mathcal{O}(1)$.
Before going further, we note briefly that a cursory glance suggests the arguments in subsections 2.1 and 3.1 should fail for $A_{0}=-2$. This is because the main scalings for the inner problem, namely (23) and (25), depend on the balance between $u_{0}$ and $tu_{1}$ (using (9)) happening when $x-x_{0}=\mathcal{O}(t^{1/3})$, whereas for $A_{0}=-2$ this balance is different (the balance between $u_{0}$ and $tu_{1}$ for $A_{0}=-2$ happens when $x-x_{0}=\mathcal{O}(t^{1/2})$). However, as we see from the higher-order terms (95)–(96), for $A_{0}=-2$ we have
@@ B.4.1 Special case $A_{0}=-2$, $A_{1}\neq 0$
In the special case $A_{0}=-2$, the leading-order problem (27)–(28) (see also (100)) has the exact solution $f_{0}=-2/\xi^{2}$ and so $\xi_{0}=0$ and we immediately see from (114) that the speed of the singularities is $\mathcal{O}(t^{-1/3})$ instead of $\mathcal{O}(t^{-2/3})$. For $A_{0}=-2$, $A_{1}\neq 0$, the problem (98), (101), becomes
$f_{1}^{\prime\prime}-\left(\frac{12}{\xi^{2}}+\xi\right)f_{1}+A_{1}=0,$ (120)
$f_{1}\sim A_{1}\left(\frac{1}{\xi}-\frac{10}{\xi^{4}}-\frac{80}{\xi^{7}}\right)\quad\mbox{as}\quad\xi\rightarrow-\mathrm{i}\infty,$ (121)
@@
In addition to (121), we are after a solution that does not grow exponentially in the lower half plane. To make that slightly more precise, we note that the homogeneous part of (120) is approximated by Airy’s equation in the far field, with linearly independent solutions
@@
As exponential growth is inconsistent with (121), we must eliminate that option. As such, we may rewrite (121) as
$f_{1}\sim A_{1}\left(\frac{1}{\xi}-\frac{10}{\xi^{4}}-\frac{80}{\xi^{7}}+\ldots\right)+\frac{K_{1}}{\xi^{1/4}}\,\mathrm{e}^{2\xi^{3/2}/3}\quad\mbox{as}\quad\xi\rightarrow-\mathrm{i}\infty,$ (122)
where the ellipsis denotes higher-order terms in a divergent expansion. Further, as with our discussion for PII in subsection 3.3, we note that, while the exponential $\xi^{-1/4}\mathrm{e}^{2\xi^{3/2}/3}$ decays as $\xi\rightarrow-\mathrm{i}\infty$, it becomes of $\mathcal{O}(1)$ on the anti-Stokes line $\theta\sim-\pi/3$ and becomes exponentially large for $\theta>-\pi/3$. Therefore, in order to exclude any exponential growth in the lower half $\xi$ plane, we also supplement (120)–(122) with
$K_{1}=0.$ (123)
With this in mind, a more complete description of the problem for $f_{1}$ is (120), (122)–(123).
In practice, it is easiest to apply (122) along the anti-Stokes line $\theta=-\pi/3$ instead of the negative imaginary axis, as setting $K_{1}=0$ then kills off both exponentials, as required. Thus we can change variables via
@@
$g_{1}^{\prime\prime}+\left(-\frac{12}{\eta^{2}}+\eta\right)g_{1}=A_{1},$ (124)
$g_{1}\sim A_{1}\left(\frac{1}{\eta}+\frac{10}{\eta^{4}}-\frac{80}{\eta^{7}}+\ldots\right)+\frac{K_{1}}{\eta^{1/4}}\,\mathrm{e}^{-2\mathrm{i}\eta^{3/2}/3+\pi\mathrm{i}/12}\quad\mbox{as}\quad\eta\rightarrow+\infty\quad\mbox{with}\quad K_{1}=0.$ (125)
The ode (124) is an inhomogeneous version of Bessel’s equation, with exact solution
@@
Thus, by using the result $\xi_{1}=-\mu_{1}/4$ (from (113)), we conclude the initial double pole at $x=x_{0}$ moves as
$s(t)\sim x_{0}+(3t)^{2/3}\frac{2\pi A_{1}\mathrm{e}^{\pi\mathrm{i}/3}}{3^{7/6}\Gamma({\textstyle\frac{2}{3}})}\quad\mbox{as}\quad t\rightarrow 0^{+}.$ (126)
@@ B.4.2 Special case $A_{0}=-2$, $A_{1}=0$, $A_{2}\neq 0$
An even more special case, which happens to be relevant for the 1-soliton solution (or, indeed, if we take any existing solution of KdV (1) and interpret it as an initial condition), is when $A_{0}=-2$ but $A_{1}=0$. Here, we have $f_{0}=-2/\xi^{2}$ ($\xi_{0}=0$) and $f_{1}=0$ ($\xi_{1}=0$), and we need to go to the next order to solve for $f_{2}$. With these solutions for $f_{0}$ and $f_{1}$, (99) and (102) become
$\frac{24}{\xi^{3}}f_{2}-\left(\frac{12}{\xi^{2}}+\xi\right)f_{2}^{\prime}+f_{2}^{\prime\prime\prime}=0,$
@@
Thus, given that $f_{2}\sim\mu_{2}/\xi^{3}$ as $\xi\rightarrow 0$, where $\mu_{2}=-8A_{2}$, we find from the result $\xi_{2}=-\mu_{2}/4$ (in (113)) that $\xi_{2}=2A_{2}$. As a consequence, for this special case with $A_{0}=-2$, $A_{1}=0$, there is one singularity that emerges from $x_{0}$ and propagates as
$s(t)\sim x_{0}+6A_{2}t\quad\mbox{as}\quad t\rightarrow 0^{+}.$ (127)
@@ B.5 Summary
In summary, by considering higher-order terms in our asymptotic expansions, we are able to describe the small-time motion of each of the singularities of the KdV solution as (108), namely $s(t)\sim x_{0}+(3t)^{1/3}\xi_{0}+(3t)^{2/3}\xi_{1}+(3t)\xi_{2}$. Each term in (108) comes from determining the location of singularities of a cascade of problems, as follows.
  • The value $x_{0}$ represents the location of each singularity $x=x_{0}$ of the initial condition $u_{0}(x)$. Here we are mostly interested in problems for which on the real axis $u_{0}$ has no singularities and is real, whereby each of these singularities will come as a complex-conjugate pair, it being sufficient to consider only those in the upper half plane.
  • The value $\xi_{0}$ represents the location of each singularity $\xi=\xi_{0}$ of the function $f_{0}(\xi)$, which is the leading-order term for the inner expansion near $x=x_{0}$, where $\xi=(x-x_{0})/(3t)^{1/3}$. These singularities are determined as the poles of our PII problem that have a residue +1+1. Their precise location will depend on the parameter $\alpha$, which itself depends on $A_{0}$, the strength of the double pole of the initial condition. Note that generically there are infinitely many of these singularities $\xi=\xi_{0}$, which means for each $x_{0}$ in the upper half plane, there are (generically) infinitely many singularities, with (108).
@@
  • In principle, further terms in (108) could be computed; these would depend on $A_{3}$, $A_{4}$, and so on.
@@ C.1 Careful stocktake of terms in transseries
A starting point for applying our transseries expansion is to extend (62)–(63) to be
@@
$\sum_{n=0}^{\infty}\sigma_{1}^{n}\left(F^{(n)^{\prime\prime}}+2n\xi^{1/2}F^{(n)^{\prime}}+\left((n^{2}-1)\xi+{\textstyle\frac{1}{2}}n\xi^{-1/2}\right)F^{(n)}\right)\,\mathrm{e}^{2n\xi^{3/2}/3}+\alpha$
@@
$F^{(0)^{\prime\prime}}-\xi F^{(0)}+\alpha=2F^{(0)3}$
$F^{(1)^{\prime\prime}}+2\xi^{1/2}F^{(1)^{\prime}}+{\textstyle\frac{1}{2}}\xi^{-1/2}F^{(1)}=6F^{(0)2}F^{(1)}$
$F^{(2)^{\prime\prime}}+4\xi^{1/2}F^{(2)^{\prime}}+(3\xi+\xi^{-1/2})F^{(2)}=6F^{(0)2}F^{(2)}+6F^{(0)}F^{(1)2}$
$F^{(3)^{\prime\prime}}+6\xi^{1/2}F^{(3)^{\prime}}+(8\xi+{\textstyle\frac{3}{2}}\xi^{-1/2})F^{(3)}=6F^{(0)2}F^{(3)}+12F^{(0)}F^{(1)}F^{(2)}+2F^{(1)3}.$
@@
provided $F_{2m}^{(0)}=0$, $F_{0}^{(2n)}=0$, $F_{0}^{(1)}=1$. This argument justifies (65).

### v2

@@ B.3 Special case $A_{1}=0$
Very briefly, if $A_{1}=0$, then the solution for $f_{1}$ is simply $f_{1}=0$, so that (B.12) implies $\mu_{1}=0$. Further, the ode for $f_{2}$, (B.6), becomes
$\xi\frac{\mathrm{d}f_{2}}{\mathrm{d}\xi}+6\left(f_{0}\frac{\mathrm{d}f_{2}}{\mathrm{d}\xi}+\frac{\mathrm{d}f_{0}}{\mathrm{d}\xi}f_{2}\right)+\frac{\mathrm{d}^{3}f_{2}}{\mathrm{d}\xi^{3}}=0,$ (B.22)
and the local behaviour near $\xi=\xi_{0}$, namely (B.13), simplifies to
$f_{2}\sim\frac{\mu_{2}}{(\xi-\xi_{0})^{3}}\quad\mbox{as}\quad\xi\rightarrow\xi_{0}.$ (B.23)
@@
where the value $\mu_{2}$ comes from solving (B.22) subject to
$f_{2}\sim A_{2}+\frac{4A_{0}A_{2}}{\xi^{3}}+\frac{40A_{0}A_{2}(A_{0}+2)}{\xi^{6}}\quad\mbox{as}\quad\xi\rightarrow-\mathrm{i}\infty$ (B.24)
@@
An interesting check on these arguments is with $A_{0}=-6$, $A_{1}=0$ and $A_{2}=2$, which is what we have for the initial condition $u_{0}=6\,\mathrm{sech}^{2}x$ and the corresponding $2$-soliton solution (4.2). In that case, we have
@@
as can be checked by substitution into (3.5) with (3.6) and (B.22) with (B.24) using $A_{0}=-6$ and $A_{2}=2$. In this exceptional case, there are only three singularities of $f_{0}$, which are given by (3.37). For each of these, the near-singularity behaviour for $f_{2}$ is
@@
which, by comparison with (B.23), means that $\mu_{2}=4$ and the three singularities emerge out of $x=x_{0}$ according to
$s(t)\sim x_{0}+(3t)^{1/3}\xi_{0}+4(3t)\quad\mbox{as}\quad t\rightarrow 0^{+}.$ (B.25)
@@
Note the special case $A_{0}=-12$, $A_{1}=0$, $A_{2}=4$ is relevant for the initial condition $u_{0}=12\,\mathrm{sech}^{2}x$, which leads to the $3$-soliton (4.6) described in subsection 4.3. In this case, the solutions of (3.5) with (3.6) and (B.22) with (B.24) are
@@
$s(t)\sim x_{0}+(3t)^{1/3}\xi_{0}+8(3t)\quad\mbox{as}\quad t\rightarrow 0^{+}.$ (B.26)
The asymptotic prediction (B.26) agrees with the exact results for the $3$-soliton solution in subsection 4.2.
@@ B.4 Special case $A_{0}=-2$
In terms of our original problem (1.1) with (1.3), the case $A_{0}=-2$ is subtly different because this is when the initial condition has a double pole that has the same strength exhibited by the double poles that time-dependent solutions to KdV have. We therefore expect the temporal scaling to be different for $A_{0}=-2$, with any singularity moving more slowly away from $x=x_{0}$ than those that are born for $A_{0}\neq-2$.
To help clarify these cases for $A_{0}=-2$, we observe that (B.14) means that singularities of solutions of KdV are doubles poles with zero residue. Thus we have two broad cases to consider. First, if $A_{0}=-2$ but $A_{1}\neq 0$, then the initial condition has the same leading-order behaviour near $x=x_{0}$ as time-dependent solutions of KdV have near each singularity $x=s(t)$, but with a different first-order correction. It turns out, in this case, that singularities are modified at $x=x_{0}$ but move with initial speed $\mathcal{O}(t^{-1/3})$ (instead of $\mathcal{O}(t^{-2/3})$, which is what happens if $A_{0}\neq-2$). Second, if $A_{0}=-2$ and $A_{1}=0$, then the initial condition has the same leading-order and correction term near its singularity as the full time-dependent solution does. This would happen, of course, if we take any time-dependent solution of (1.1) (that has evolved for any nonzero time) and treat it as an initial condition. Here the singularity present at $x_{0}$ propagates with speed $\mathcal{O}(1)$.
Before going further, we note briefly that a cursory glance suggests the arguments in subsections 2.1 and 3.1 should fail for $A_{0}=-2$. This is because the main scalings for the inner problem, namely (3.1) and (3.3), depend on the balance between $u_{0}$ and $tu_{1}$ (using (2.3)) happening when $x-x_{0}=\mathcal{O}(t^{1/3})$, whereas for $A_{0}=-2$ this balance is different (the balance between $u_{0}$ and $tu_{1}$ for $A_{0}=-2$ happens when $x-x_{0}=\mathcal{O}(t^{1/2})$). However, as we see from the higher-order terms (B.2)–(B.3), for $A_{0}=-2$ we have
@@ B.4.1 Special case $A_{0}=-2$, $A_{1}\neq 0$
In the special case $A_{0}=-2$, the leading-order problem (3.5)–(3.6) (see also (B.7)) has the exact solution $f_{0}=-2/\xi^{2}$ and so $\xi_{0}=0$ and we immediately see from (B.21) that the speed of the singularities is $\mathcal{O}(t^{-1/3})$ instead of $\mathcal{O}(t^{-2/3})$. For $A_{0}=-2$, $A_{1}\neq 0$, the problem (B.5), (B.8), becomes
$\frac{\mathrm{d}^{2}f_{1}}{\mathrm{d}\xi^{2}}-\left(\frac{12}{\xi^{2}}+\xi\right)f_{1}+A_{1}=0,$ (B.27)
$f_{1}\sim A_{1}\left(\frac{1}{\xi}-\frac{10}{\xi^{4}}-\frac{80}{\xi^{7}}\right)\quad\mbox{as}\quad\xi\rightarrow-\mathrm{i}\infty,$ (B.28)
@@
In addition to (B.28), we are after a solution that does not grow exponentially in the lower half plane. To make that slightly more precise, we note that the homogenous part of (B.27) is approximated by Airy’s equation in the far field, with linearly independent solutions
@@
As exponential growth is inconsistent with (B.28), we must eliminate that option. As such, we may rewrite (B.28) as
$f_{1}\sim A_{1}\left(\frac{1}{\xi}-\frac{10}{\xi^{4}}-\frac{80}{\xi^{7}}+\ldots\right)+\frac{K_{1}}{\xi^{1/4}}\,\mathrm{e}^{2\xi^{3/2}/3}\quad\mbox{as}\quad\xi\rightarrow-\mathrm{i}\infty,$ (B.29)
where the ellipsis denotes higher-order terms in a divergent expansion. Further, as with our discussion for PII in subsection 3.4, we note that, while the exponential $\xi^{-1/4}\mathrm{e}^{2\xi^{3/2}/3}$ decays as $\xi\rightarrow-\mathrm{i}\infty$, it becomes of $\mathcal{O}(1)$ on the anti-Stokes line $\theta=-\pi/3$ and becomes exponentially large for $\theta>-\pi/3$. Therefore, in order to exclude any exponential growth in the lower half $\xi$-plane, we also supplement (B.27)–(B.29) with
$K_{1}=0.$ (B.30)
With this in mind, a more complete description of the problem for $f_{1}$ is (B.27), (B.29)–(B.30).
In practice, it is easiest to apply (B.29) along the anti-Stokes line $\theta=-\pi/2$ instead of the negative imaginary axis, as setting $K_{1}=0$ then kills off both exponentials, as required. Thus we can change variables via
@@
$\frac{\mathrm{d}^{2}g_{1}}{\mathrm{d}\xi^{2}}+\left(-\frac{12}{\eta^{2}}+\eta\right)g_{1}=A_{1},$ (B.31)
$g_{1}\sim A_{1}\left(\frac{1}{\eta}+\frac{10}{\eta^{4}}-\frac{80}{\eta^{7}}+\ldots\right)+\frac{K_{1}}{\eta^{1/4}}\,\mathrm{e}^{-2\mathrm{i}\eta^{3/2}/3+\pi\mathrm{i}/12}\quad\mbox{as}\quad\eta\rightarrow+\infty\quad\mbox{with}\quad K_{1}=0.$ (B.32)
The ode (B.31) is an inhomogeneous version of Bessel’s equation, with exact solution
@@
Thus, by using the result $\xi_{1}=-\mu_{1}/4$ (from (B.20)), we conclude the initial double pole at $x=x_{0}$ moves as
$s(t)\sim x_{0}+(3t)^{2/3}\frac{2\pi A_{1}\mathrm{e}^{\pi\mathrm{i}/3}}{3^{7/6}\Gamma({\textstyle\frac{2}{3}})}\quad\mbox{as}\quad t\rightarrow 0^{+}.$ (B.33)
@@ B.4.2 Special case $A_{0}=-2$, $A_{1}=0$, $A_{2}\neq 0$
An even more special case, which happens to be relevant for the 1-soliton solution (or, indeed, if we take any existing solution of KdV (1.1) and interpret it as an initial condition), is when $A_{0}=-2$ but $A_{1}=0$. Here, we have $f_{0}=-2/\xi^{2}$ ($\xi_{0}=0$) and $f_{1}=0$ ($\xi_{1}=0$), and we need to go to the next order to solve for $f_{2}$. With these solutions for $f_{0}$ and $f_{1}$, (B.6) and (B.9) become
$\frac{24}{\xi^{3}}f_{2}-\left(\frac{12}{\xi^{2}}+\xi\right)\frac{\mathrm{d}f_{2}}{\mathrm{d}\xi}+\frac{\mathrm{d}^{3}f_{2}}{\mathrm{d}\xi^{3}}=0,$
@@
Thus, given that $f_{2}\sim\mu_{2}/\xi^{3}$ as $\xi\rightarrow 0$, where $\mu_{2}=-8A_{2}$, we find from the result $\xi_{2}=-\mu_{2}/4$ (in (B.20)) that $\xi_{2}=2A_{2}$. As a consequence, for this special case with $A_{0}=-2$, $A_{1}=0$, there is one singularity that emerges from $x_{0}$ and propagates as
$s(t)\sim x_{0}+6A_{2}t\quad\mbox{as}\quad t\rightarrow 0^{+}.$ (B.34)
@@ B.5 Summary
In summary, by considering higher-order terms in our asymptotic expansions, we are able to describe the small-time motion of each of the singularities of the KdV solution as (B.15), namely $s(t)\sim x_{0}+(3t)^{1/3}\xi_{0}+(3t)^{2/3}\xi_{1}+(3t)\xi_{2}$. Each term in (B.15) comes from determining the location of singularities of a cascade of problem as follows:
The value $x_{0}$ represents the location of each singularity $x=x_{0}$ of the initial condition $u_{0}(x)$. Here we are mostly interested in problems for which $u_{0}$ is real on the real axis, whereby each of these singularities will come as a complex-conjugate pair, it being sufficient to consider only those in the upper half plane.
The value $\xi_{0}$ represents the location of each singularity $\xi=\xi_{0}$ of the function $f_{0}(\xi)$, which is the leading-order term for the inner expansion near $x=x_{0}$, where $\xi=(x-x_{0})/(3t)^{1/3}$. These singularities are determined as the poles of our PII problem that have a residue +1+1. Their precise location will depend on the parameter $\alpha$, which itself depends on $A_{0}$, the strength of the double pole of the initial condition. Note that generically there are infinitely many of these singularities $\xi=\xi_{0}$, which means for each $x_{0}$ in the upper half plane, there are (generically) infinitely many singularities, with (B.15).
@@
In principle, further terms in (B.15) could be computed; these would depend on $A_{3}$, $A_{4}$, and so on.
@@ C.1 Careful stocktake of terms in transseries
A starting point for applying our transseries expansion is to extend (3.43)–(3.44) to be
@@
$\sum_{n=0}^{\infty}\sigma_{1}^{n}\left(\frac{\mathrm{d}^{2}F^{(n)}}{\mathrm{d}\xi^{2}}+2n\xi^{1/2}\frac{\mathrm{d}F^{(n)}}{\mathrm{d}\xi}+\left((n^{2}-1)\xi+{\textstyle\frac{1}{2}}n\xi^{-1/2}\right)F^{(n)}\right)\,\mathrm{e}^{2n\xi^{3/2}/3}+\alpha$
@@
$\frac{\mathrm{d}^{2}F^{(0)}}{\mathrm{d}\xi^{2}}-\xi F^{(0)}+\alpha=2(F^{(0)})^{3}$
$\frac{\mathrm{d}^{2}F^{(1)}}{\mathrm{d}\xi^{2}}+2\xi^{1/2}\frac{\mathrm{d}F^{(1)}}{\mathrm{d}\xi}+{\textstyle\frac{1}{2}}\xi^{-1/2}F^{(1)}=6(F^{(0)})^{2}F^{(1)}$
$\frac{\mathrm{d}^{2}F^{(2)}}{\mathrm{d}\xi^{2}}+4\xi^{1/2}\frac{\mathrm{d}F^{(2)}}{\mathrm{d}\xi}+(3\xi+\xi^{-1/2})F^{(2)}=6(F^{(0)})^{2}F^{(2)}+6F^{(0)}(F^{(1)})^{2}$
$\frac{\mathrm{d}^{2}F^{(3)}}{\mathrm{d}\xi^{2}}+6\xi^{1/2}\frac{\mathrm{d}F^{(3)}}{\mathrm{d}\xi}+(8\xi+{\textstyle\frac{3}{2}}\xi^{-1/2})F^{(3)}=6(F^{(0)})^{2}F^{(3)}+12F^{(0)}F^{(1)}F^{(2)}+2(F^{(1)})^{3}.$
@@
provided $F_{2m}^{(0)}=0$, $F_{0}^{(2n)}=0$, $F_{0}^{(1)}=1$. This argument justifies (3.46).